# Strang splitting estimator for nonlinear multivariate stochastic differential equations with Pearson-type multiplicative noise

### 17 Apr 2026 v1

@@ Abstract
Multivariate Pearson diffusions, also known as polynomial diffusions, are characterized by a linear drift vector and a diffusion matrix that is quadratic in the state variables. We derive exact closed-form expressions for the mean and covariance matrix of this class by using results on matrix exponential integrals. We then extend this framework to a broader class of nonlinear diffusions with Pearson-type multiplicative noise. The main contribution of this paper is a new parameter estimator for these nonlinear models based on Strang splitting (SS). The proposed method decomposes the stochastic system into a deterministic nonlinear ordinary differential equation (ODE) and a multivariate Pearson diffusion. We construct the SS estimator by composing their respective flows and applying a Gaussian transition approximation parameterized by the exact moments of the Pearson component. We prove that the SS estimator is consistent and asymptotically efficient. Furthermore, we introduce a new model within this broader class, which we call the Student Kramers oscillator, and we prove existence and uniqueness of the strong solution as well as existence of an invariant measure. We evaluate the SS estimator through simulation studies on this new oscillator and on the multivariate Wright-Fisher diffusion from population genetics. These simulations demonstrate that the SS estimator outperforms the standard Euler-Maruyama estimator, the Gaussian approximation estimator, and the local linearization estimator. Finally, we apply the SS estimator to fit the Student Kramers oscillator to Greenland ice core data.
@@ A Pearson diffusion is the solution to a one-dimensional SDE of the form (Forman and Sørensen, 2008)
[…]
@@ 5 Application to Greenland Ice Core Data
We consider three nested models of increasing complexity, summarised in Table 1. The models $\mathrm{M}_{1}^{[\text{old}]}$ and $\mathrm{M}_{1}$ are identical in structure, with $b=d=\alpha=\beta=0$, corresponding to a symmetric double-well potential with constant diffusion. $\mathrm{M}_{1}^{[\text{old}]}$ uses the parameter estimates from (Pilipovic et al., 2025), while $\mathrm{M}_{1}$ re-estimates the same model using our method. The negative log-likelihood is lower for $\mathrm{M}_{1}$, which shows that the estimates in (Pilipovic et al., 2025) are not optimal. The intermediate model $\mathrm{M}_{2}$ retains Student-type noise by freeing $\alpha$ and $\beta$, but keeps $b=d=0$, so the potential $U(x)$ remains symmetric. The full model $\mathrm{M}_{3}$ frees all eight parameters.
@@
The model fit is assessed in the right panels of Figure 6. We present the empirical invariant distributions of both coordinates alongside the fitted theoretical invariant distribution. A $95\%$ prediction interval is constructed by simulating 1000 datasets from the fitted model matching the size of the empirical data, and taking pointwise 2.5th and 97.5th percentiles of the resulting approximated densities. A single example trajectory is shown in dark red. Note that the theoretical invariant densities are approximated under the assumption that $X_{t}$ and $V_{t}$ are decoupled.
Compared to the standard Kramers oscillator $\mathrm{M}_{1}$, the Student Kramers oscillator $\mathrm{M}_{2}$ improves the fit for the velocity variable $V_{t}$, whose heavy-tailed empirical distribution is better captured by the Student-type noise. However, the fit for $X_{t}$ becomes worse since the estimated stable equilibria are closer together than the observations, leading to the model placing substantial probability mass between the two states. Thus, the implied transitions are smoother and more gradual than the nearly discontinuous jumps seen in the data.
To further compare the models to data, we estimate the mean occupancy times in each climate state via successive up- and down-crossings of the smoothed process at predefined thresholds, following the same procedure as in (Pilipovic et al., 2025), with one correction. Namely, occupancy times are obtained by multiplying crossing counts by the observation interval of 20 years, rather than dividing them by $h=0.02$ as reported there. The resulting mean occupancy times are shown in Table 2, and confirm the improvement of $\mathrm{M}_{2}$ over $\mathrm{M}_{1}$. While $\mathrm{M}_{1}$ underestimates the mean persistence in both states at 771 and 762 years, $\mathrm{M}_{2}$ yields 913 and 844 years, much closer to the observed 1172 and 841 years.
Table 2: Estimated mean occupancy times (in years) in stadial and interstadial states.
Stadial (years) Interstadial (years)
Data 1172 841
$\mathrm{M}_{1}$ (standard Kramers) 771 762
$\mathrm{M}_{2}$ (Student Kramers) 913 844
The underestimation of the stadial duration, together with the inability to reproduce abrupt transitions, suggests that the model is still too simple to fully capture the dynamics, and that more complex models may be necessary to account for the abrupt nature of the Dansgaard–Oeschger transitions.

### 23 Apr 2026 v2

@@ Abstract
Multivariate Pearson diffusions are characterized by a linear drift and a diffusion matrix that is quadratic in the state variables. We derive closed-form expressions for the mean and covariance matrix of this class using matrix exponential integrals, and extend this framework to a broader class of nonlinear diffusions with Pearson-type multiplicative noise. The main contribution is a new parameter estimator for these nonlinear multiplicative models based on Strang splitting, which decomposes the stochastic system into a deterministic nonlinear ordinary differential equation and a multivariate Pearson diffusion. The estimator is constructed by composing their respective flows and applying a Gaussian transition approximation with exact moments from the Pearson component. We prove that the estimator is consistent and asymptotically efficient. We also introduce a new model within this class, the Student Kramers oscillator, and prove existence and uniqueness of the strong solution and of an invariant measure. We evaluate the estimator through simulation studies on this oscillator and on the multivariate Wright-Fisher diffusion from population genetics, where it outperforms the Euler-Maruyama, Gaussian approximation, and local linearization estimators. We conclude with an application to Greenland ice core data using the Student Kramers oscillator.
@@ A Pearson diffusion is the solution to a one-dimensional SDE of the form (Forman and Sørensen, 2008)
Notation. Capital bold letters denote random vectors, vector-valued functions, and matrices, lowercase bold letters denote deterministic vectors. $\|\cdot\|$ denotes the $L^{2}$ vector norm in $\mathbb{R}^{d}$. Superscript $(i)$ on a vector denotes the $i$-th component. The double subscript $ij$ on a matrix denotes the component in the $i$-th row and $j$-th column. $\operatorname*{Tr}(\cdot)$ returns the trace, $\det(\cdot)$ the determinant, $\top$ the transpose, $\operatorname*{vec}$ vectorization, and $\operatorname*{diag}(\cdot)$ either the diagonal matrix formed from a vector or the vector of diagonal entries of a matrix. The Kronecker product and the sum of two matrices are $\otimes$ and $\oplus$, respectively. $\mathbf{I}_{d}$ denotes the $d$-dimensional identity matrix, $\bm{0}_{d\times d}$ a $d\times d$-dimensional zero matrix, $\mathbf{1}$ a vector of ones. We denote by $a_{i}]_{i=1}^{d}$ a vector with coordinates $a_{i}$, and by $b_{ij}]_{i,j=1}^{d}$ a matrix with coordinates $b_{ij}$. For a function $\bm{g}$, $\partial_{x^{(i)}}\bm{g}(\mathbf{x})$ denotes the partial derivative with respect to $x^{(i)}$ and $\partial_{x^{(i)}x^{(j)}}^{2}\bm{g}(\mathbf{x})$ with respect to $x^{(i)}$ and $x^{(j)}$. The nabla operator $\nabla_{\mathbf{x}}$ denotes the gradient vector of real-valued $g$ with respect to $\mathbf{x}$; $\nabla_{\mathbf{x}}g(\mathbf{x})=[\partial_{x^{(i)}}g(\mathbf{x})]_{i=1}^{d}$. For function $\mathbf{F}:\mathbb{R}^{d}\to\mathbb{R}^{d}$, $D$ denotes the Jacobian matrix $D\mathbf{F}(\mathbf{x})=[\partial_{x^{(i)}}F^{(j)}(\mathbf{x})]_{i,j=1}^{d}$. $\mathbf{R}$ represents a vector (or matrix) valued function defined on $(0,1)\times\mathbb{R}^{d}$ (or $(0,1)\times\mathbb{R}^{d\times d}$), such that, for some constant $C$, $\|\mathbf{R}(a,\mathbf{x})\|<aC(1+\|\mathbf{x}\|)^{C}$ for all $a,\mathbf{x}$; $R$ refers to a scalar function. $\overline{A}$ denotes the closure of an open set $A$. $\delta_{ij}$ denotes the Kronecker delta. We write $\mathbf{X}\sim\mathcal{N}_{d}(\bm{\mu},\bm{\Omega})$ for $\mathbf{X}$ a $d$-dimensional Gaussian with mean $\bm{\mu}$ and covariance $\bm{\Omega}$; in general, $\sim$ denotes distribution. $\propto$ denotes proportionality up to a constant, and $\approx$ denotes a numerical approximation. We write $\xrightarrow{\mathbb{P}},\xrightarrow{d},\xrightarrow{a.s.}$ for convergence in probability, in distribution, and almost surely, respectively.
@@ 5 Application to Greenland Ice Core Data
We consider three nested models of increasing complexity, summarised in Table 1. $\mathrm{M}_{1}$, with $b=d=\alpha=\beta=0$, corresponds to a symmetric double-well potential with constant diffusion. The negative log-likelihood for $\mathrm{M}_{1}$ shows that the estimates in (Pilipovic et al., 2025) ($\widehat{\eta}=62.5$, $\widehat{a}=-219$, $\widehat{d}=297$, $\widehat{\gamma}=9125$, and $\text{NLL}=8708.05$) are not optimal. The intermediate model $\mathrm{M}_{2}$ retains Student-type noise by freeing $\alpha$ and $\beta$, but keeps $b=d=0$, so the potential $U(x)$ is symmetric. The full model $\mathrm{M}_{3}$ frees all eight parameters.
@@
The model fit is assessed in the right panels of Figure 6. We present the empirical invariant distributions of both coordinates alongside the fitted theoretical invariant distribution. A $95\%$ prediction interval is constructed by simulating 1000 datasets from the fitted model matching the size of the empirical data, and taking pointwise 2.5th and 97.5th percentiles of the resulting approximated densities. A single example trajectory is shown in red. Note that the theoretical invariant densities are approximated under the assumption that $X_{t}$ and $V_{t}$ are decoupled.
Compared to the standard Kramers oscillator $\mathrm{M}_{1}$, the Student Kramers oscillator $\mathrm{M}_{2}$ improves the fit for the heavy-tailed velocity variable $V_{t}$. However, the fit for $X_{t}$ becomes worse since the estimated stable equilibria are closer together than the observations, leading to the model placing substantial probability mass between the two states. Thus, the implied transitions are smoother and more gradual than the nearly discontinuous jumps seen in the data.
The inability to reproduce abrupt transitions, suggests that the model is still too simple to fully capture the dynamics, and that more complex models may be necessary to account for the abrupt nature of the Dansgaard–Oeschger transitions.